LA-MB-FTMW STUDIES OF SUGARS

M. LOZOYA, C. CABEZAS, S. MATA, J. C. LÓPEZ, \underline{J. L. ALONSO}, Grupo de Espectroscopía Molecular (GEM). Edificio Quifima. Laboratorios de Espectroscopía y Bioespectroscopía. Parque Científico. Universidad de Valladolid, 47011 Valladolid. (Spain).

Recently, rotational studies of biomolecules have entered in a new stage with the LA-MB-FTMW experiment.^{a} It combines laser ablation with Fourier transform microwave spectroscopy in supersonic jets overcoming the problems of thermal decomposition associated with conventional heating methods. This technique has been successfully applied to the study of monosaccarides. Three conformers of the prototype α-D-glucose and other three for β-D-glucose have been characterized for the first time in the gas phase.

aJ. L. Alonso, C. Pérez, M. E. Sanz, J. C. López, S. Blanco, Phys. Chem. Chem. Phys. 11,617-627 (2009)and references therein